# PRING: Rethinking Protein-Protein Interaction Prediction from Pairs to Graphs

### v1

@@ 1 Introduction
To fill this blank, we propose PRING, a multi-organisms holistic benchmark that evaluates the PRotein-protein INteraction prediction from a Graph perspective. As illustrated in Fig. 1, PRING compiles high-confidence physical interactions across four organisms, Human, Arath, Ecoli, and Yeast, sourced from STRING szklarczyk2023string , UniProt uniprot2023uniprot , Reactome milacic2024reactome , and IntAct del2022intact , yielding 21,484 proteins and 186,818 interactions, with efforts made to minimize both data redundancy and leakage. Through this golden dataset, we benchmark representative PPI prediction methods, including sequence similarity-based, naive sequence-based, PLM-based models, and structure-based, across five tasks: two topology-oriented, namely intra-species and cross-species PPI network construction, to assess whether models can recover the topology properties of PPI networks, such as network density and local community structures, and facilitating cross-species biological knowledge transfer study, and three function-oriented, including protein complex pathway prediction, GO functional module analysis, and essential protein justification, which can support disease mechanism analysis, protein function annotation, and therapeutic target identification. Collectively, these evaluations reveal whether a model is capable of capturing both the topological and the functional semantics of real interactomes, moving beyond pairwise classification to network-level understanding.
Extensive experiments yield the following insights: (1) current models tend to generate overly dense graphs, diverging from the sparsity nature of real PPI networks; (2) predicted PPI modules exhibit limited functional alignment with the ground-truth, restricting their utility in downstream tasks such as pathway reconstruction and function annotation; and (3) reconstructed graphs struggle to separate essential from non-essential proteins, indicating that critical topological signals remain uncaptured; (4) classification metrics cannot completely reflect a model’s ability to recover network structure, highlighting the need for graph-level evaluations. While these findings highlight challenges, we hope they can serve as a positive step forward for the community. PRING complements existing benchmarks and provides a reliable platform for more holistic research in modeling PPI networks.
@@ 3.1 Data Collection
Resource Integration. We begin by aggregating experimentally validated PPIs from UniProt uniprot2023uniprot , Reactome milacic2024reactome , and IntAct del2022intact , and complement these sources with high-confidence interactions (combined score > 0.7 islam2023molecular ) from STRING szklarczyk2023string to ensure both both data quality and comprehensive coverage, while mitigating the noise inherent in some STRING interactions. To further enhance data reliability and prepare for downstream function-oriented analyses, we retain only annotated proteins with known functions in SwissProt bairoch2000swiss . Additionally, four phylogenetically distinct organisms, including Human, Arabidopsis thaliana (Arath), Escherichia coli (Ecoli), and Yeast, are selected using NCBI Taxonomy identifiers schoch2020ncbi to match with the target species and increase the species diversity. This step not only considers data quality but also emphasizes data diversity and coverage, providing a critical foundation for constructing our gold-standard dataset.
@@ 3.2.1 Topology-Oriented Task
To evaluate whether a PPI prediction model can faithfully recover the topology of interaction networks, we first formulate the PPI network reconstruction task as follows. Given a PPI graph with $N$ proteins, each protein is represented by its features $x_{i}=\{s_{i},c_{i}\},\quad i=1,\dots,N$, where $s_{i}$ is the amino-acid sequence and $c_{i}$ is any auxiliary context. The model $f$ predicts an interaction label for each protein pair $(i,j)$ with $1\leq i<j\leq N$, according to: $I^{\mathrm{pred}}_{i,j}=f(x_{i},x_{j})\in\{0,1\}$. Here, $I^{\mathrm{pred}}_{i,j}=1$ indicates a predicted interaction between protein $i$ and protein $j$, while $I^{\mathrm{pred}}_{i,j}=0$ denotes no predicted interaction. We then aggregate all positive predictions to form the reconstructed network:
[…]
@@ 4 Experiment
We conduct extensive experiments on five network-level tasks outlined in Sec. 3.2. Our experiments aim to answer the following research questions: Q1: How effectively do current models reconstruct PPI networks? (Sec. 4.3.1) Q2: How well do the PPI prediction models generalize? (Sec. 4.3.2) Q3: How do these models perform in biology practice? (Sec. 4.3.3) Q4: How well do standard classification metrics reflect network-level performance? (Sec. 4.3.4)
@@ 4.1 Evaluation Baseline
Structure-based Model. These models either take known protein structures as input or use end-to-end structure prediction for PPI modeling. We include three structure-based approaches: Struct2Graph baranwal2022struct2graph and TAGPPI song2022learning , which use GNNs to model protein structure input, and RF2-Lite humphreys2024protein , a lightweight variant of RoseTTAFold2 baek2023efficient designed specifically for PPI prediction. In addition, a case study of Chia-1 chai2024chai , the newest structure prediction model, is provided in Apx. G.3.
4.2 Experimental Setup
@@
Naive Sequence Topsy-Turvy 0.183 1.74 17.3 11.4 14.6 0.130 3.22 180 201 65.9 0.232 4.88 104 64.8 36.7 5
@@ 4.3.2 Beyond Boundaries: Generalization Across Species
The second topology-oriented task evaluates the cross-species generalization ability of PPI prediction models. As shown in Fig. 3, the overall performance consistently decreases as the evolutionary distance from Human increases, following the order of Arath, Yeast, and Ecoli, with the average performance drops by 15.2%, 25.3%, and 35.2%, respectively. In terms of model categories, PLM-based models continue to achieve the best performance among all baselines, with an average graph similarity score of 0.24 across the three species. This result indicates the robustness and generalizability of PLM-derived representations across diverse biological scenarios. In contrast, naive sequence-based models perform the worst, with an average score of only 0.17, demonstrating that simple deep learning architectures such as CNNs and RNNs are insufficient for capturing the complex evolutionary and structural signals required for cross-species.
@@
Naive Sequence Topsy-Turvy 0.240 0.613 0.338 0.232 0.635 0.43 0.755 8
@@ A.1 PRING Dataset
The PRING dataset is available at https://doi.org/10.7910/DVN/22AUPR.
@@ C.1 Data Sources
In this work, we construct the PPI benchmark dataset using data integrated from multiple public data sources, including UniProt, IntAct, Reactome, STRING, Gene Ontology (GO), and the Consensus CDS (CCDS) project. License information for each source is provided in Tab. 4.
@@ C.2 Collection Process
CC   -!- INTERACTION:
CC       P12345; Q8N1H7:XYZ_HUMAN (xeno); Xeno interaction [EMBL:ABC123];
CC       P12345; Q9Y6K1:DEF_HUMAN; NbExp=3; IntAct=EBI-12345, EBI-67890;
If the interacting partners is a Swiss-Prot protein, we extract the interaction as a PPI entry. Additionally, we download the full Swiss-Prot FASTA file containing 573,230 protein sequences. Using the NCBI taxonomy ID specified in the FASTA headers (e.g., OX=9606), we extract protein sets for four representative species: Human (Homo sapiens, 9606), Arath (Arabidopsis thaliana, 3702), Yeast (Saccharomyces cerevisiae, 559292), and Ecoli (Escherichia coli, 83333).
@@ C.3 Statistical Analysis
We conduct a statistical analysis of the fully processed PPI datasets obtained through the above pipeline, with results summarized as follows:
@@
Tab. 5 summarizes key statistics of the processed PPI datasets across the four selected species. The number of proteins and interactions varies notably after the filter process, with Human (contributing the training set) having the largest network (129,861 interactions among 10,090 proteins), while Yeast and Ecoli have more compact networks. In terms of the topology, Human exhibits the smallest diameter and average path length, suggesting a denser and more interconnected PPI network. In contrast, Arath shows a larger diameter and average shortest path length, indicating a more fragmented or modular interaction structure.
@@ D.1 Topology-oriented Task
Here, $I^{\mathrm{pred}}_{i,j}=1$ indicates a predicted interaction, and 00 indicates no interaction.
@@
[…]
@@ D.1.1 Comparison of Different Traversal Strategies
BFS Locally dense structures, such as protein complexes, modules Ribosomal protein complex, intracellular signaling module [101]
DFS Chain/tree structures, such as signaling or metabolic pathways MAPK signaling pathway, glycolysis metabolic pathway [102]
RW Hub-dominant structures, such as regulatory networks and pan-functional networks TP53 regulatory network, transcription factor network [103]
@@ D.1.4 Evaluation Metrics
Graph $\text{Graph Similarity}=1-\frac{\|\hat{A}-A\|_{1}}{|E|+|\hat{E}|},$ (4)
where $A\in\{0,1\}^{N\times N}$ is the adjacency matrix of the ground-truth graph $G^{\text{true}}$, $\hat{A}\in\{0,1\}^{N\times N}$ is the adjacency matrix of the predicted graph $G^{\text{pred}}$. $E$ and $\hat{E}$ are the sets of edges in the ground-truth and predicted graphs, respectively.
@@
Relative $\text{Relative Density}=\frac{|\hat{E}|/\binom{N}{2}}{|E|/\binom{N}{2}}=\frac{|\hat{E}|}{|E|}$ (5)
where $N$ is the number of proteins in the PPI network, and $|E|$ and $|\hat{E}|$ denote the number of edges in the ground-truth and predicted graphs, respectively.
@@
[…]
@@ D.2.1 Protein Complex Pathway Prediction
Pathway $\text{Pathway Precision}=\frac{|E^{\text{pred}}\cap E^{\text{true}}|}{|E^{\text{pred}}|},$ (7)
@@
Pathway $\text{Pathway Recall}=\frac{|E^{\text{pred}}\cap E^{\text{true}}|}{|E^{\text{true}}|},$ (8)
@@
Pathway $\text{Pathway Connectivity}=\frac{1}{N}\sum_{i=1}^{N}\left[G^{\text{pred}}_{i}\text{ is connected}\right],$ (9)
@@ D.2.2 GO Enrichment Analysis
Functional $\text{Functional Alignment}=\frac{1}{K}\sum_{i=1}^{K}\max_{j}\frac{|T_{i}^{\text{pred}}\cap T_{j}^{\text{true}}|}{|T_{i}^{\text{pred}}\cup T_{j}^{\text{true}}|},$ (10)

### v2

@@ 1 Introduction
To fill this blank, we propose PRING, a multi-organisms holistic benchmark that evaluates the PRotein-protein INteraction prediction from a Graph perspective. As illustrated in Fig. 1, PRING compiles high-confidence physical interactions across four organisms, Human, Arath, Ecoli, and Yeast, sourced from STRING szklarczyk2023string , UniProt uniprot2023uniprot , Reactome milacic2024reactome , and IntAct del2022intact , yielding 21,484 proteins and 186,818 interactions, with efforts made to minimize both data redundancy and leakage. Through this golden dataset, we benchmark representative PPI prediction methods, including sequence similarity-based, naive sequence-based, PLM-based models, and structure-based, across five tasks: two topology-oriented, namely intra-species and cross-species PPI network construction, to assess whether models can recover the topology properties of PPI networks, such as network density and local community structures, and facilitating cross-species biological knowledge transfer study, and three function-oriented, including protein complex pathway prediction, GO functional module analysis, and essential protein justification, which can support disease mechanism analysis, protein function annotation, and therapeutic target identification. Collectively, these evaluations reveal whether a model is capable of capturing both the topological and the functional semantics of real interactomes, moving beyond pairwise classification to network‑level understanding.
Extensive experiments yield the following insights: (1) current models tend to generate overly dense graphs, diverging from the sparsity nature of real PPI networks; (2) predicted PPI modules exhibit limited functional alignment with the ground-truth, restricting their utility in downstream tasks such as pathway reconstruction and function annotation; and (3) reconstructed graphs struggle to separate essential from non‑essential proteins, indicating that critical topological signals remain uncaptured; (4) classification metrics cannot completely reflect a model’s ability to recover network structure, highlighting the need for graph-level evaluations. While these findings highlight challenges, we hope they can serve as a positive step forward for the community. PRING complements existing benchmarks and provides a reliable platform for more holistic research in modeling PPI networks.
@@ 2.2 Protein-Protein Interaction Prediction Model
2.3 Multi-Modal Learning for Enhanced Biological Representation
An emerging research direction seeks to overcome the limitations of single-modality PPI models by constructing richer representations from multimodal data. One of the approaches is to unify molecular information, such as structure and sequence, with semantic data like natural language to improve functional prediction. This paradigm is exemplified by models that connect protein data to text for better interpretability liu2024prott3 ; zhang2023ontoprotein ; gao2023biotranslator . The approach has been broadly explored by fusing chemical structures with text liu2023molca ; 3dmolm ; simsgt ; liu2024reactxt ; liu2025nextmol ; luo2025towards . Such advancements in unified learning offer a promising path to developing more powerful PPI models that can address the topological and functional challenges identified by the PRING benchmark.
@@ 3.1 Data Collection
Resource Integration. We begin by aggregating experimentally validated PPIs from UniProt uniprot2023uniprot , Reactome milacic2024reactome , and IntAct del2022intact , and complement these sources with high-confidence interactions (combined score > 0.7 islam2023molecular ) from STRING szklarczyk2023string to ensure both data quality and comprehensive coverage, while mitigating the noise inherent in some STRING interactions. To further enhance data reliability and prepare for downstream function-oriented analyses, we retain only annotated proteins with known functions in SwissProt bairoch2000swiss . Additionally, four phylogenetically distinct organisms, including Human, Arabidopsis thaliana (Arath), Escherichia coli (Ecoli), and Yeast, are selected using NCBI Taxonomy identifiers schoch2020ncbi to match with the target species and increase the species diversity. This protein filter step not only considers data quality but also emphasizes data diversity and coverage, providing a critical foundation for constructing our gold-standard dataset.
@@ 3.2.1 Topology-Oriented Task
To evaluate whether a PPI prediction model can faithfully recover the topology of interaction networks, we first formulate the PPI network reconstruction task as follows. Given a PPI graph with $N$ proteins, each protein is represented by its features $x_{i}=\{s_{i},c_{i}\},\quad i=1,\dots,N$, where $s_{i}$ is the amino‐acid sequence and $c_{i}$ is any auxiliary context. The model $f$ predicts an interaction label for each protein pair $(i,j)$ with $1\leq i<j\leq N$, according to: $I^{\mathrm{pred}}_{i,j}=f(x_{i},x_{j})\in\{0,1\}$. Here, $I^{\mathrm{pred}}_{i,j}=1$ indicates a predicted interaction between protein $i$ and protein $j$, while $I^{\mathrm{pred}}_{i,j}=0$ denotes no predicted interaction. We then aggregate all positive predictions to form the reconstructed network:
$G^{\mathrm{pred}}=\mathrm{Aggregate}\bigl\{(i,j)\mid I^{\mathrm{pred}}_{i,j}=1,\;1\leq i<j\leq N\bigr\}.$ (1)
@@ 4 Experiment
We conduct extensive experiments on five network-level tasks outlined in Sec. 3.2. Our experiments aim to answer the following research questions: Q1: How effectively do current models reconstruct PPI networks? (Sec. 4.3.1) Q2: How well do the PPI prediction models generalize? (Sec. 4.3.2) Q3: How do these models perform in biology practice? (Sec. 4.3.3) Q4: How well do standard classification metrics reflect network-level performance? (Sec. 4.3.4) Q5: How does the positive–negative ratio affect model performance and network reconstruction? (Sec. 4.3.5)
@@ 4.1 Evaluation Baseline
Structure-based Model. These models either take known protein structures as input or use end-to-end structure prediction for PPI modeling. We include three structure-based methods: Struct2Graph baranwal2022struct2graph and TAGPPI song2022learning , which use GNNs to model protein structure input, and RF2-Lite humphreys2024protein , a lightweight variant of RoseTTAFold2 baek2023efficient designed specifically for PPI prediction. In addition, a case study of Chai-1 chai2024chai , the newest structure prediction model, is provided in Apx. G.5.
@@
Naive
Sequence
Topsy-Turvy 0.183 1.74 17.3 11.4 14.6 0.130 3.22 180 201 65.9 0.232 4.88 104 64.8 36.7 5
@@ 4.3.2 Beyond Boundaries: Generalization Across Species
The second topology-oriented task evaluates the cross-species generalization ability of PPI prediction models. As shown in Fig. 3, the overall performance decreases proportionately as the evolutionary distance from Human increases, following the order of Arath, Yeast, and Ecoli, with the average performance drops by 15.2%, 25.3%, and 35.2%, respectively. In terms of model categories, PLM-based models continue to achieve the best performance among all baselines, with an average graph similarity score of 0.24 across the three species. This result indicates the robustness and generalizability of PLM-derived representations across diverse biological scenarios. In contrast, naive sequence-based models perform the worst, with an average score of only 0.17, demonstrating that simple deep learning architectures such as CNNs and RNNs are insufficient for capturing the complex evolutionary and structural signals required for cross-species.
@@
Naive
Sequence
Topsy-Turvy 0.240 0.613 0.338 0.232 0.635 0.43 0.755 8
@@ 4.3.4 Metric Alignment: Do Classification Scores Tell the Whole Story?
Figure 6: Ablation on the ratio of positive to negative PPI pairs. Results are shown for the BFS-based sampling configuration. For relative density, a value closer to 1 denotes better topological consistency.
4.3.5 Data Matters: Do Positive to Negative Ratios Affect Network Construction?
Considering that real PPI networks are inherently sparse, we vary the positive-to-negative ratio of training pairs (1:1 to 1:10) under the BFS-based sampling configuration to assess the effect of class imbalance (Fig. 6). Since SPRINT does not require negative samples for training and the PLM-interact models are fixed for inference, we include four representative methods in this analysis. The complete results across three sampling strategies are provided in Apx. G.3.
For D-SCRIPT and Struct2Graph, increasing the ratio severely degrades performance—the models tend to underpredict interactions, with RD scores approaching 0. In contrast, PPITrans and TUnA show improved results when the ratio increases to 1:5, suggesting moderate imbalance regularizes learning, but their performance drops again at 1:10 as excessive negatives bias the networks toward overly sparse topologies.
Findings. Moderate class imbalance can enhance the performance of stronger models such as PLM-based architectures by providing richer negative supervision and better reflecting the sparsity of real PPI networks. However, the optimal positive-to-negative ratio remains task- and architecture-dependent, warranting further investigation into more advanced strategies for biologically realistic network learning.
@@ 5 Conclusion
Acknowledgment
This work was partially supported by the New Generation Artificial Intelligence-National Science and Technology Major Project of China (2025ZD0121801). This work was also partially supported by the Ministry of Education, Singapore (MOE T1251RES2309, MOE T2EP20125-0039) and the Agency for Science, Technology and Research (A*STAR H25J6a0034). This work was also partially supported by the JC STEM Lab of AI for Science and Engineering, funded by The Hong Kong Jockey Club Charities Trust, the Research Grants Council of Hong Kong (Project No. CUHK14213224). This work was also partially supported by Shanghai Artificial Intelligence Laboratory.
@@ A.1 PRING Dataset
The PRING dataset is available at https://huggingface.co/datasets/piaolaidangqu/PRING.
@@ A.2 Evaluation Code
A.3 Project Website
The PRING project website is available at https://zhanglab.comp.nus.edu.sg/PRING/.
@@ C.1 Data Sources
In this work, we construct the PPI benchmark dataset using data integrated from multiple public data sources. For PRING PPI network, we select UniProt, Reactome, IntAct, and STRING as our primary source databases. These databases are chosen based on the following considerations: 1) They collectively provide high-quality annotations and broad biological coverage; 2) Most other PPI databases either overlap substantially with these sources (e.g., BioGrid [109], MINT [110]) or exhibit limited scope and lower data reliability (e.g., HuRI [111], PrePPI [112]); 3) These four resources are widely recognized and commonly adopted within the bioinformatics community [113, 114].
For function-oriented tasks, the data includes Gene Ontology (GO), and the Consensus CDS (CCDS) project. License information for each source is provided in Tab. 4.
@@ C.2 Collection Process
$C$   -!- INTERACTION:
$C$       P12345; Q8N1H7:XYZ_HUMAN (xeno); Xeno interaction [EMBL:ABC123];
$C$       P12345; Q9Y6K1:DEF_HUMAN; NbExp=3; IntAct=EBI-12345, EBI-67890;
If the interacting partners are a Swiss-Prot protein, we extract the interaction as a PPI entry. Additionally, we download the full Swiss-Prot FASTA file containing 573,230 protein sequences. Using the NCBI taxonomy ID specified in the FASTA headers (e.g., OX=9606), we extract protein sets for four representative species: Human (Homo sapiens, 9606), Arath (Arabidopsis thaliana, 3702), Yeast (Saccharomyces cerevisiae, 559292), and Ecoli (Escherichia coli, 83333).
@@ C.3 Statistical Analysis
To ensure transparency in dataset construction, we first report the detailed preprocessing steps for building the PPI benchmark. After retrieving interaction data from four curated databases, we first apply a protein-level filter (Fig. 2) to retain only experimentally verified proteins with annotated functions (SwissProt) and belonging to one of the four target species. This filtering step results in 15,043 Human, 13,271 Arath, 5,232 Yeast, and 3,855 E. coli proteins.
We then perform sequence similarity filtering ($\leq$40%) using MMSeqs2, followed by function-based redundancy removal to eliminate homologous or functionally overlapping proteins across species.
  • Arath: 7,487 proteins were removed by MMSeqs2, and 759 were excluded due to shared functions with Human, yielding 5,025 final proteins (↓62%).
  • Yeast: 1,033 proteins were filtered by MMSeqs2, and 1,026 shared functions with Human, leaving 3,173 proteins (↓39%).
  • Ecoli: 452 proteins were removed by MMSeqs2, and 207 due to function overlap, resulting in 3,196 proteins (↓17%).
After these filtering steps, we obtain the final PRING PPI network dataset.
Then, we conduct a statistical analysis of the fully processed PPI datasets obtained through the above pipeline, with results summarized as follows:
@@
Tab. 5 summarizes key statistics of the processed PPI datasets across the four selected species. The number of proteins and interactions varies notably after the filter process, with Human (contributing to the training set) having the largest network (129,861 interactions among 10,090 proteins), while Yeast and Ecoli have more compact networks. In terms of the topology, Human exhibits the smallest diameter and average path length, suggesting a denser and more interconnected PPI network. In contrast, Arath shows a larger diameter and average shortest path length, indicating a more fragmented or modular interaction structure.
@@ D.1 Topology-oriented Task
Here, $I^{\mathrm{pred}}_{i,j}=1$ indicates a predicted interaction, and 0 indicates no interaction.
@@
$G^{\mathrm{pred}}=\mathrm{Aggregate}\bigl\{(i,j)\mid I^{\mathrm{pred}}_{i,j}=1,\;1\leq i<j\leq N\bigr\}.$ (3)
@@ D.1.1 Comparison of Different Traversal Strategies
BFS
Locally dense structures, such as
protein complexes, modules
Ribosomal protein complex, intra-
cellular signaling module
[115]
DFS
Chain/tree structures, such as
signaling or metabolic pathways
MAPK signaling pathway, glycoly-
sis metabolic pathway
[116]
RW
Hub-dominant structures, such as
regulatory networks and pan-
functional networks
TP53 regulatory network, transcrip-
tion factor network
[117]
@@ D.1.4 Evaluation Metrics
$\text{Graph Similarity}=1-\frac{\|\hat{A}-A\|_{1}}{|E|+|\hat{E}|}$, (4)
[…]
@@
$\text{Relative Density}=\frac{|\hat{E}|/\binom{N}{2}}{|E|/\binom{N}{2}}=\frac{|\hat{E}|}{|E|}$ (5)
where $N$ is the number of proteins in the PPI network, $and |E^||\hat{E}|$ denote the number of edges in the ground-truth and predicted graphs, respectively.
@@
$\text{Degree Distribution (MMD)}=\frac{\text{MMD}^{2}(\text{pred},\text{test})}{\text{MMD}^{2}(\text{test},\text{test})},$ (6)
@@ D.2.1 Protein Complex Pathway Prediction
$\text{Pathway Precision}=\frac{|E^{\text{pred}}\cap E^{\text{true}}|}{|E^{\text{pred}}|},$ (7)
@@
$\text{Pathway Recall}=\frac{|E^{\text{pred}}\cap E^{\text{true}}|}{|E^{\text{true}}|},$ (8)
@@
$\text{Pathway Connectivity}=\frac{1}{N}\sum_{i=1}^{N}\left[G^{\text{pred}}_{i}\text{ is connected}\right],$ (9)
@@ D.2.2 GO Enrichment Analysis
$\text{Functional Alignment}=\frac{1}{K}\sum_{i=1}^{K}\max_{j}\frac{|T_{i}^{\text{pred}}\cap T_{j}^{\text{true}}|}{|T_{i}^{\text{pred}}\cup T_{j}^{\text{true}}|},$ (10)